### v1

# Galvanometer-scanning transient phase microscopy with balanced detection and arbitrary pump polarization

Galvanometer-scanning transient phase microscopy with balanced detection and arbitrary pump polarization
††journal: opticajournal††articletype: Research ArticleTransient absorption microscopy measures excited-state kinetics based on the imaginary part of the pump-induced perturbation to the complex refractive index, i.e. , with applications in both materials and biomedical sciences. Its complement, transient phase microscopy, enabled by stable inline birefringent interferometry, measures the real part . The ability to switch between absorption and phase measurements may yield a stronger signal, depending on the sample and probe wavelength. To date, however, transient phase has not been coupled with galvanometer scanners, thus limiting it to materials science applications and non-imaging spectroscopy. Here, we extend transient phase microscopy to operate in a fast galvanometer-scanning microscope with balanced detection, comparing amplitude and phase measurements in graphene, hemoglobin, and red blood cells. We examine the impacts and limitations introduced by galvanometer scanning, in addition to relocation of the pump-probe combining dichroic to permit arbitrary polarization of the pump.
1 Introduction
Pump-probe transient absorption spectroscopy is a valuable tool for probing femtosecond chemical reactions and excited-state molecular relaxation pathways. [Zewail_1988] When coupled with a laser-scanning microscope, transient absorption microscopy (TAM) [Fischer_Wilson_Robles_Warren_2016] generates contrast from absorptive materials through pump-induced changes in the complex refractive index (e.g. excited-state absorption, ground state depletion, stimulated emission, transient changes in Franck-Condon broadening, and transient photoproducts). TAM can also measure, at each pixel, lifetimes associated with relaxation pathways and kinetics (e.g. spontaneous emission, vibrational cooling, and photodissociated ligand recombination). The differences in excited-state lifetimes can provide contrast between materials that have very similar absorption spectra, such as melanins [MelaninPP_1_Warren_2015, MelaninPP_2_Warren_2016, MelaninPP_3_Warren_2019]. Other applications include hemoglobin spectroscopy and imaging [HemoglobinPP_1_Fu_2019, HemoglobinPP_2_Fu_2020] and emerging techniques to evaluate mitochondrial redox through cytochrome electron transport hemeproteins [ErkangPP_1_Wilson_2022, ErkanPP_2_Wilson_2022], along with numerous applications in materials sciences [MaterialSciApp_1_Li_2021, MaterialSciApp_2_Wang_2019, MaterialSciApp_3_Ma_2022, ExtractingQuantitativeDielectricFxn_2022] and cultural heritage [HeritagePP_1_Fischer_2014, HeritagePP_2_Ji_2023].
With few exceptions, TAM operates by measuring pump-induced changes to the detected probe intensity of a beam that has traversed a sample, following excitation by a pump pulse with a time delay of i.e. , where denotes the imaginary part of a complex quantity. The complementary measurement by transient phase microscopy (TM), senses , where denotes the real part, through interference between the probe and a reference pulse, although diffraction, scattering, or lensing could also be exploited. [bartels2021low] In pump-probe spectroscopy, such phase measurements are most often associated with impulsive Raman scattering (ISRS), where molecular bond deformations or rotations are directly coupled to via the differential polarizability [bartels2001phase, bartels2002nonresonant, Wilson_JosaB_THGVibSpec_2012, bartels2021low] and can be measured through interferometry [hartinger2008single, Wilson_SyntheticAperture_2008, wilson2008phase, schlup2009sensitive, MichelsonInterferometer_Wahlstrand_2023] or spectral shifting [domingue2014time, Smith_IRS_Bartels_2021, smith2022nearly]. A similar situation is found with pump-induced molecular rotational motion [Robles_PPNLDS_Warren_2013]. On the other hand, in pump-probe imaging and spectroscopy of absorptive pigments, is largely neglected, owing to the simplicity of measuring , which does not require stable probe-reference interference. However, full characterization of requires both real and imaginary parts, and there may be some scenarios where TM has advantages over TAM. For example, consider the ground state bleaching response of an absorption band. After pumping to the excited state, the maximal change in absorptivity will occur at the peak wavelength of the ground state absorption band. Thus, for an absorption measurement, the probe wavelength yielding the strongest response will be at the absorption peak, and will thus achieve poor penetration into thick materials. For a phase measurement, however, the maximum change in refractive index response inferred from Kramers-Kronig relations will be offset from the absorption peak. Thus, the probe wavelength yielding the strongest phase response will not suffer the same depth limitations.
Various techniques have been developed to measure through its modulation of probe pulse phase. The simplest measurement detects spectral shifting proportional to the time derivative of the modulation, as has been done in cross-phase modulation (XPM) imaging of biological samples [XPMSS_wilson] and is frequently employed in ISRS spectroscopy and imaging [bartels2021low, Smith_IRS_Bartels_2021]. A more direct measurement of probe phase can be done by interference with a reference pulse. Early work in ISRS focused on spectral interferometry, where the time-delayed probe and reference pulses produce interference fringes on a spectrometer [SagnacInterferometer_Kobayashi_1995, MichelsonInterferometer_Wahlstrand_2023]. The fringe spacing is inversely proportional to probe-reference delay, and pump-induced causes the fringes to shift. Alternatively, probe and reference pulses can be re-timed and overlapped on a single-element detector [van2005detection, Orrit, Guillet_2015]. Such measurements are spoiled by any relative phase instability between probe and reference, and high stability is achieved in geometries where probe and reference pulses trace a common path from the point where they are split from a common initial pulse to the point where their interference is measured after interaction with the sample.
The challenge in common-path interferometry is to ensure that the probe interacts with the pump-induced perturbations to the sample, while the reference does not. This can be done in a Sagnac interferometer in which counter-propagating probe and reference pulses arrive at different times in the sample [SagnacInterferometer_Kobayashi_1995], or by propagating probe and reference along slightly different paths so that the pump overlaps spatially only with the probe [MichelsonInterferometer_Wahlstrand_2023], or by using group delay differences in bulk birefringent media to create a probe-reference pair, which arrive at different times in the sample [Schlup]. (It should be noted that for short probe-reference delays, these methods measure a finite difference rather than directly [Wilson_SyntheticAperture_2008, bartels2021low].) Of these, the birefringent common path interferometer provides a simple means of re-timing probe and reference, enabling measurement with a single-element detector rather than a spectrometer, and has been employed for microscopic spectroscopy. In a transmission setup with a pair of pulse-splitting and re-timing calcites, [Orrit] measured both the real and imaginary components of the perturbed optical transmission coefficient for gold nanoparticles. Similarly, [Guillet_2015] made complex measurements of the perturbed optical reflection coefficient for a tungsten thin film. In their setup, the probe and reference are split on their way to the sample by a birefringent crystal, then after reflection from the sample re-timed by propagating in the reverse direction through the same crystal. A half-wave plate can shift the measurement from phase to amplitude in both studies, permitting measurement of the complex . This approach to TM has yet to be implemented on a galvanometer-scanning microscope, and thus has not been applied to imaging of biological samples. In addition, the placement of the pump-combining dichroic before the birefringent splitter precludes rotation of pump polarization angle, which can have a large effect on time-resolved spectroscopy of biological pigments such as melanin [Grass_PolarizationMelaninPP_Warren_2022].
Here, we extend TM with birefringent inline interferometry to a galvanometer-scanning microscope, measuring biological samples in a transmission geometry. In contrast to previous works [Orrit, Guillet_2015], we place the pump-combining dichroic after the probe-reference birefringent splitter, allowing for the freedom to arbitrarily rotate the pump polarization. We evaluate the impact that galvanometer scanning has on probe-reference re-timing and field of view, in addition to the signal-to-noise impact of polarization effects from placing the dichroic after the birefringent probe-reference splitter [DichroicAberrations]. Furthermore, we implement a balanced detection scheme that doubles the TM signal while canceling relative intensity noise from the laser. We demonstrate imaging applications on graphene and red blood cells (using two-photon absorption for pumping), comparing the relative signal-to-noise ratio (SNR) achievable between TAM and TM. Our results demonstrate the feasibility of galvanometer-scan imaging using transient phase detection and suggest that the ability to switch between TAM and TM can yield an improved SNR.
2 Methods
This research presents a comparison of time-resolved transient absorption versus time-resolved transient phase measurements. We utilize a pair of birefringent calcite crystals for separation and re-timing of the probe and reference pulses, then measure the temporal interference of the probe-reference pair following the pump-sample interaction.
2.1 Experimental Setup
Our laser source is generated by a Ytterbium-doped fiber laser oscillator (Menlo YLMO) which seeds a fiber chirped pulse amplifier that provides pulses at centered at with a bandwidth of nm. This output is split into a pump pulse train for the experiments, and another portion feeds a noncollinear optical parametric amplifier to generate the probe and reference pulses at with a bandwidth of nm. Afterwards, dispersion compensation is managed on the pump arm with a grating compressor and a prism compressor, and on the probe arm with a prism compressor.
The system diagram after the laser source is shown in Fig. 1. The probe arm is directed through delay stage which controls the delay between the pump and probe pulse trains. A set of polarization optics (HWP 1, POL 1, HWP 2) cleans up the probe polarization and rotates it to with respect to the eigenaxes of a 2 mm thick calcite (CAL 1), with its optic axis oriented in the plane of the table. This results in an equal projection of the probe pulse along the fast (extraordinary) and slow (ordinary) axes of the crystal, respectively, resulting in two output pulses. The reference pulse exits first and is p-polarized followed by a delayed probe pulse that is s-polarized (see Supplementary 1A). The pump arm is directed through an acousto-optic modulator (AOM), which applies a 500 kHz sinusoidal amplitude modulation. A polarizer and half-wave plate provides rotatable pump polarization, before combining with the probe-reference pair in the dichroic mirror (DM) (Thorlabs DMLP900). The combined pulses are directed into a laser scanning microscope, with a pair of galvanometernometer scan mirrors conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm.
A pump reject filter eliminates the pump pulse train after the condenser and the reference-probe pulse pair is relayed from the condenser to the re-timing calcite (CAL 2). CAL 2 is oriented so as to undo the relative delay between the probe and reference, with its slow (ordinary) axis aligned with the advanced reference pulse and fast (extraordinary) axis aligned with the delayed probe pulse. At the exit of the re-timing calcite, probe and reference pulses are temporally overlapped and in quadrature ( relative offset phase, adjusted by fine-tuning the azimuthal angle of the splitter calcite) but orthogonally polarized. A half-wave plate then rotates probe and reference so they are at with respect to the s- and p- polarization axes of a polarizing beam splitter (PBS). Detector A (Thorlabs PDA36A2) measures p-polarized light transmitted through the PBS and Detector B (Thorlabs PDA36A2) measures s-polarized light reflected from the PBS. In this configuration, an advance in probe phase relative to the pump increases A while decreasing B. By rotating the half-wave plate, detection can be reconfigured for TAM, rather than TM, as detailed below. Throughout, relay lenses are used to image the Fourier plane (collection objective back aperture) to the re-timing calcite and the detectors, minimizing the effects of lateral movement during image scanning applications.
2.2 Detection Technique
A diagram focusing on the polarization optics is shown in Figure 2. The probe beam proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1). The polarizer ensures that we have linearly polarized light at step 1 which is p-polarized (parallel to the plane of the optics table). HWP 2 is used to rotate the linearly polarized light to 45 degrees (step 2). CAL 1 is a 2 mm thick piece of calcite cut so that the optic axis runs along the front face of the crystal (a-cut). It is oriented so that the extraordinary axis coincides with p-polarized light and the ordinary axis coincides with s-polarized light. When the probe beam is incident on CAL 1 at 45 degrees it is projected onto the ordinary () and extraordinary axes (). Here and for are the refractive index and group index, evaluated at the center frequency of the probe pulse, . Two pulses exit CAL 1 (step 3) orthogonally polarized; the reference pulse is advanced and p-polarized, the probe pulse is delayed and s-polarized. The probe pulse pair proceeds through the microscope before encountering CAL 2. CAL 2 is oriented opposite to CAL 1; the ordinary (slow) axis is rotated to coincide with the advanced reference pulse and the extraordinary (fast) axis coincides with the delayed probe pulse. The result is a re-timing of the two pulses (step 4). Note that the axes of CAL 2 may not align with s- and p-polarization as referenced to the table surface, but is instead based on the axes of the reference and probe polarizations after going through the microscope. HWP 3 is used to rotate the re-timed reference and probe polarizations such that they are 45 degrees to the s- and p- axes of the polarizing beam splitter (PBS), so that both the reference and probe pulses send a copy of themselves to DET A and DET B (step 5). The pulse pair copy on DET B (step 6) will be out of phase from the pulse pair on DET A (step 7). Thus, if one detector measures destructive interference between the reference-probe pair then the other detector will measure constructive interference.
With each sample, the initial offset phase () between the probe and reference pulses is set to a balanced position such that , where is the voltage on DET A and is the voltage on DET B. The offset phase can be set by adjusting the angle-of-incidence () of CAL 1 with p-polarized light. This angle is set with the pump pulse blocked so that the offset phase accounts for any residual phase imparted to the reference and probe pulses by the microscope and sample. When the pump pulse is unblocked, a phase advance imparted to the probe pulse from an excited sample will cause an increase in and a decrease in and the overall balanced detection signal will swing positive. Similarly, a phase delay imparted to the probe pulse will cause a decrease in and an increase in and the overall signal will swing negative. Figure 3 shows an example of the temporal interference between the reference (pink) and probe (blue) pulses on detector A and B for these scenarios.
Ultimately our setup is capable of measuring an excited-state phase change of from the initial offset point. The voltage measured on each detector is the time-integrated interference between the reference and probe; since our final signal is we end up with a 2X increase in measurable signal than if we were conducting our measurements with only one detector. The preamplifier (SR 560) measures and applies a voltage gain before the difference signal is sent into a lock-in amplifier (Liquid Instruments Moku:Lab). The lock-in amplifier measures amplitude changes of that coincide with the 500 kHz modulation frequency of the pump train; these measurements correspond to pump-induced phase changes .
The balanced detection signal is given by
where is a phase offset between reference and probe set by CAL1, CAL2, any birefringent optics in the path, e.g. the pump-combining dichroic, and static birefringence of the sample (see Supplementary 2 and 3 for more details). The term arising from pump-probe interactions is . In the balanced configuration, , and assuming that the phase perturbations resulting from are small,
One advantage of this detection configuration is the ability to switch between measuring transient absorption () and transient phase () simply by changing the projection of the probe and reference pulses on the PBS axes [Orrit, Guillet_2015]. This projection angle is controlled by HWP 3; Figure 4 shows three projection scenarios of the reference-probe pair polarization axes onto the PBS. If HWP 3 is oriented such that the reference-probe pair polarization axes are 45 degrees to the s- and p- axes of the PBS (left image), then there is an equal amplitude projection of both pulses to the two detectors. This setting is for making purely a phase measurement since any changes to the amplitudes of either pulse will cancel in the balanced detection scheme; only changes to the temporal interference on either detector caused by changes in will result in .
If HWP 3 is oriented such that the reference-probe pair polarization axes are aligned to the s- and p- axes of the PBS (right image), then DET A will measure amplitude changes to the reference pulse and DET B will measure amplitude changes to the probe pulse. This setting is for making purely an absorption measurement since any changes to the phase of either pulse will have no bearing on the integrated intensity measured by the detector; only changes to the amplitude of either pulse caused by changes in will result in .
Finally, if HWP 3 is oriented for any other projection (middle image), then a combination of a phase and absorption measurement will be made.
3 Results
3.1 Transient Phase and Absorption Measurements
This section presents the results of measuring transient absorption and transient phase in glass, a graphene monolayer, bovine hemoglobin in solution and in live human red blood cells. The probe wavelength of 800 nm is considered off-resonance for the hemoglobin and red blood cell samples.
3.1.1 Cross-Phase Modulation Measurement in Glass
First we acquired a delay scan in a glass microscope slide (Figure 5). Glass has a minimal absorption at the probe wavelength of 800 nm and the interaction that occurs between the sample and the pump pulse is instantaneous. Thus, the width of the XPM pulses gives us an estimate of pulse duration for our reference and probe pulses as they overlap with the pump pulse in time; these pulses are 350 fs in duration.
3.1.2 Graphene Monolayer
Next we explored the 2D and 3D nonlinear response of a graphene monolayer on a 0.5 mm quartz substrate placed on top of a 1 mm glass microscope slide. Previous works looking at graphene have utilized transient absorption to analyze graphene carrier dynamics [GrapheneTAS_SingleLayerHyperspectral_Gesuele_2019, MaterialSciApp_3_Ma_2022]. It is a material that shows promise for use in various applications including electronics and drug delivery, making it a popular research target [Graphene_TAM_InBlood_Yang_2015]. Here, graphene serves as a control material to compare TAM with TM since it has a strong transient absorption response at our pump and probe wavelengths.
Time-Domain Spectroscopy
An initial delay scan taken in the center of the sample revealed a stronger transient absorption versus transient phase response with a decay time of 2 picoseconds.
The first peak at ps corresponds to the pump-reference overlap and the second peak at ps corresponds to the pump-probe overlap.
Imaging
Image stacks were acquired at the edge of the graphene sample where we expected to see some variation in the deposition of graphene on the quartz substrate. A probe transmissivity image was acquired after setting the half-waveplate for phase measurements and the angle of incidence for the first calcite for a "balanced" state (i.e. ) with the pump blocked. From Figure 7 the edge of the sample can be seen and gives context to interpreting the image stack in Figure 8.
An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared with output from the lock-in amplifier (V) (Figure 8). The transient absorption measurements still provide better SNR, but similar structural information can be seen in both techniques.
3.1.3 Hemoglobin and Red Blood Cells
Next we moved onto biological samples. First, we looked at the spectroscopy of a hemoglobin sample prepared with 9 mM of bovine hemoglobin lyophilized powder (Sigma-Aldrich H2500) in phosphate buffer solution. Then we imaged a droplet of whole blood placed on a poly-L-lysine coated microscope slide with a glass coverslip. The whole blood used was acquired from healthy donors by venipuncture by Dr. Liszt Madruga of the Department of Chemical Engineering (CSU) working under Prof. Matt Kipper of the Departments of Chemical & Biological Engineering (CSU) and Prof. Ketul Popat of the Department of Mechanical Engineering (CSU). The protocol followed was approved by the Colorado State University Institutional Review Board.
Time-Domain Spectroscopy of Hemoglobin
The hemoglobin sample showed a decay curve (top row in Figure 9). Interestingly, hemoglobin demonstrated a pump polarization-dependent signal with differences between the polarization sensitivity of the transient absorption versus transient phase measurements (bottom row of Figure 9). This would be impossible to observe if the pump-combining dichroic was placed before the probe-reference splitting calcite like in [Orrit, Guillet_2015]. With transient absorption the pump-reference overlap showed a polarization dependence while the pump-probe overlap did not. For the pump-reference overlap, the strongest signal occurred with an s-polarized pump. With transient phase both overlaps demonstrated a polarization dependence; the pump-reference overlap shows the largest signal for the s-polarized pump and the pump-probe overlap shows the largest signal for the p-polarized pump.
Red Blood Cell Imaging
Imaging human red blood cells, we expected to see a predominant signal from hemoglobin and a similar SNR from the two techniques (based on the results from the delay scans). However, the transient phase demonstrated a stronger signal than transient absorption in this case with clearer structural details due to XPM. Figure 10 shows a comparison of TM and TAM for imaging the transient decay in red blood cells. All of the images are shown on the same intensity scale; the overlaps with the pump-reference at ps and with the pump-probe at ps show the strongest signals and good structural detail for TM.
4 Discussion
4.1 Impacts of Dichroic Beam-Combiner on SNR
The pump-combining dichroic was placed after the probe-reference splitting calcite, unlike prior works which placed the dichroic before the splitter to avoid splitting the pump [Orrit, Guillet_2015]. This enabled characterization of polarization dependence, as well as selection of relative pump-probe polarization that maximizes signal for imaging. It should be noted, however, that the dielectric coating of the dichroic is subject to polarization aberrations, including dichroism (polarization-dependent absorption), depolarization, dispersion, and retardance [DichroicAberrations]. These effects can depend on both incident polarization state and angle of incidence, and can ultimately interfere with our detection scheme. One behavior we observed with our setup was a polarization-dependent spectral phase difference imparted to the probe and reference from the dichroic; the pulse broadening effects were minimal and inconsequential with our pulse durations (>300 fs), though we did notice a variation in spectral fringe visibility as a function of wavelength and the development of ellipticity on the probe and reference pulses. Additionally, our reference and probe pulses were found to have unequal amplitudes at the detectors. For TAM, only the amplitude differences contribute to a reduced SNR due to a DC offset. For TM, all three effects contribute to a reduced SNR. Thus, future work to address these issues will likely further the advantage of TM over TAM in this scenario.
4.2 Impact of Galvanometer-scanning on Re-timing Phase in Imaging
Galvanometer-scanning microscopy is a new development in TM and comes at the price of an angle-dependent re-timing phase present across the scanned field of view. In order to determine how great an impact galvanometer-dependent re-timing phase has on this system, a cross-phase modulation (XPM) image was taken in a glass slide at different fields of view. Since glass is homogeneous and has no transient signal, a flat signal would be expected across the entire field of view in the absence of galvanometer-dependent re-timing phase. The measured data shown in Figure 11 indicates a phase shift that depends on scan angle. The images shown in Figure 11 have had the on-axis phase shift DC subtracted so only the spatially dependent phase shift is portrayed.
Clearly, the larger field of view sees a stronger galvanometer-dependent phase shift . Overall, the signal measured across a field of view will equal the sum + . Decreasing the field of view or increasing the magnification of the beam at the re-timing calcite (CAL 2) could help minimize the re-timing phase effects (Supplementary 4). In order to completely eliminate this effect, de-scanning could be added in prior to CAL 2.
4.3 Trade-offs
One significant benefit of measuring TM is that the instantaneous cross-phase modulation (XPM) is measured in addition to the transient phase signal. The XPM contributes additional context about the structure of the sample related to nonlinear refractive index variations across the sample [XPMSS_wilson]. TAM will also measure instantaneous phenomena, such as two-photon absorption, but this does not give the same structural clarity as XPM and is instead specific only to materials and molecules resonant with the sum of the pump and probe energies.
The microscope design presented here allows for a simple conversion between transient phase and transient absorption measurements. However, it utilizes an orthogonally polarized probe-reference pair. The probe-reference pair is required to perform the time-integrated interferometric phase measurements, but it is not required to make transient absorption measurements. Indeed, often TAM is conducted with only a pump-probe pulse pair. One downside of this design is that any sample phenomenon or optical element that spoils the polarization relationship of the probe-reference pair impacts the overall detected signal. This was discussed above for the case of the dichroic beam splitter, but the effects of birefringent samples can also be considered.
For thin, weakly birefringent samples the polarization relationship of the reference and probe pulses is at risk of changing, and we would expect the effects to be similar to the dichroic in reducing SNR. Thick birefringent samples run the risk of not only changing the polarization relationship of the reference and probe pulses but also of additional pulse splitting or temporal separation. Depending on the degree to which this occurs, a very small amount of the exiting reference and probe might be within the re-timing range of the secondary calcite, or the pulses may not be able to be re-timed by the calcite at all. This would preclude the ability of this system to make phase measurements entirely.
5 Conclusion
The purpose of this research was to extend TM to a galvanometer-scanning configuration, as a complement to TAM for biological samples. In sum, the scanning galvanometer TM approach described here can offer advantages over TAM, such as XPM structural contrast and potentially higher signal levels, but is susceptible to degradation from polarization aberrations in the optical setup and sample birefringence, and has more limited field of view. Thus, it is beneficial to be able to acquire both TAM and TM with a simple rotation of a waveplate, as described here. In the future two changes to the microscope design here could be made to improve the utility for TM. Firstly, the dichroic beam combiner could be swapped for a 50/50 beamsplitter. While this would cut the power at the combiner by half, it would also eliminate the polarization-dependent effects imparted by the dichroic that we have shown negatively impact the SNR of TM. Secondly, de-scanning could be added prior to the re-timing calcite to remove the spatial dependence on re-timing phase. Overall, TM shows promise as a complementary technique to TAM, particularly for setups that may be limited in their ability to vary probe wavelengths.
Funding This research was supported by the National Institute of General Medical Sciences of the National Institutes of Health under award number R21GM135772 and by the Biophotonics Program of the National Science Foundation under award number 1943595.
Acknowledgments
Disclosures The authors declare no conflicts of interest.
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Supplemental document See Supplement 1 for supporting content.

### v2

@@
 To date, however, transient phase has not been coupled with galvanometer scanners, thus limiting it to materials science applications and non-imaging spectroscopy.
-Here, we extend transient phase microscopy to operate in a fast galvanometer-scanning microscope with balanced detection, comparing amplitude and phase measurements in graphene, hemoglobin, and red blood cells.
+Here, we extend transient phase microscopy to operate in a galvanometer-scanning microscope with balanced detection, comparing amplitude and phase measurements in graphene (in which amplitude detection has the advantage), hemoglobin and red blood cells (in which phase detection has the advantage).
 We examine the impacts and limitations introduced by galvanometer scanning, in addition to relocation of the pump-probe combining dichroic to permit arbitrary polarization of the pump.
 1 Introduction
-Pump-probe transient absorption spectroscopy is a valuable tool for probing femtosecond chemical reactions and excited-state molecular relaxation pathways.
-[Zewail_1988] When coupled with a laser-scanning microscope, transient absorption microscopy (TAM) [Fischer_Wilson_Robles_Warren_2016] generates contrast from absorptive materials through pump-induced changes in the complex refractive index (e.g.
+Pump-probe transient absorption spectroscopy is a valuable tool for probing femtosecond chemical reactions and excited-state molecular relaxation pathways [1].
+When coupled with a laser-scanning microscope, transient absorption microscopy (TAM) [2] generates contrast from absorptive materials through pump-induced changes in the complex refractive index (e.g.
 excited-state absorption, ground state depletion, stimulated emission, transient changes in Franck-Condon broadening, and transient photoproducts).
@@
 spontaneous emission, vibrational cooling, and photodissociated ligand recombination).
-The differences in excited-state lifetimes can provide contrast between materials that have very similar absorption spectra, such as melanins [MelaninPP_1_Warren_2015, MelaninPP_2_Warren_2016, MelaninPP_3_Warren_2019].
-Other applications include hemoglobin spectroscopy and imaging [HemoglobinPP_1_Fu_2019, HemoglobinPP_2_Fu_2020] and emerging techniques to evaluate mitochondrial redox through cytochrome electron transport hemeproteins [ErkangPP_1_Wilson_2022, ErkanPP_2_Wilson_2022], along with numerous applications in materials sciences [MaterialSciApp_1_Li_2021, MaterialSciApp_2_Wang_2019, MaterialSciApp_3_Ma_2022, ExtractingQuantitativeDielectricFxn_2022] and cultural heritage [HeritagePP_1_Fischer_2014, HeritagePP_2_Ji_2023].
-With few exceptions, TAM operates by measuring pump-induced changes to the detected probe intensity of a beam that has traversed a sample, following excitation by a pump pulse with a time delay of i.e.
+The differences in excited-state lifetimes can provide contrast between materials that have very similar absorption spectra, such as melanins [3, 4, 5].
+Other applications include hemoglobin spectroscopy and imaging [6, 7] and emerging techniques to evaluate mitochondrial redox through cytochrome electron transport hemeproteins [8, 9], along with numerous applications in materials sciences [10, 11, 12, 13] and cultural heritage [14, 15].
+With few exceptions, TAM operates by measuring pump-induced changes to the detected amplitude of a probe pulse, following excitation after a delay of , i.e.
 , where denotes the imaginary part of a complex quantity.
-The complementary measurement by transient phase microscopy (TM), senses , where denotes the real part, through interference between the probe and a reference pulse, although diffraction, scattering, or lensing could also be exploited.
-[bartels2021low] In pump-probe spectroscopy, such phase measurements are most often associated with impulsive Raman scattering (ISRS), where molecular bond deformations or rotations are directly coupled to via the differential polarizability [bartels2001phase, bartels2002nonresonant, Wilson_JosaB_THGVibSpec_2012, bartels2021low] and can be measured through interferometry [hartinger2008single, Wilson_SyntheticAperture_2008, wilson2008phase, schlup2009sensitive, MichelsonInterferometer_Wahlstrand_2023] or spectral shifting [domingue2014time, Smith_IRS_Bartels_2021, smith2022nearly].
-A similar situation is found with pump-induced molecular rotational motion [Robles_PPNLDS_Warren_2013].
+In these experiments, images are formed by raster scanning the combined, focused, pump and probe beams across the sample.
+This can be done either by mechanically translating the sample (typical for materials science applications), or by imaging a pair of galvanometer-mounted mirrors to the back aperture of a microscope objective to scan the beam (typical for biomedical applications).
+Galvanometer scanning is more complex, but has the advantage of speed, enabling the capture of dynamics and reducing sample heating by reducing pixel dwell time.
+The complementary measurement by transient phase microscopy (TM), senses , where denotes the real part, through interference between the probe and a reference pulse, although diffraction, scattering, or lensing could also be exploited [16].
+In pump-probe spectroscopy, such phase measurements are most often associated with impulsive Raman scattering (ISRS), where molecular bond deformations or rotations are directly coupled to via the differential polarizability [17, 18, 19, 16] and can be measured through interferometry [20, 21, 22, 23, 24] or spectral shifting [25, 26, 27].
+A similar situation is found with pump-induced molecular rotational motion [28].
 On the other hand, in pump-probe imaging and spectroscopy of absorptive pigments, is largely neglected, owing to the simplicity of measuring , which does not require stable probe-reference interference.
-However, full characterization of requires both real and imaginary parts, and there may be some scenarios where TM has advantages over TAM.
+While similar in their ability to measure transient kinetics, there may be some scenarios where TM has advantages over TAM.
 For example, consider the ground state bleaching response of an absorption band.
@@
 Thus, the probe wavelength yielding the strongest phase response will not suffer the same depth limitations.
-Various techniques have been developed to measure through its modulation of probe pulse phase.
-The simplest measurement detects spectral shifting proportional to the time derivative of the modulation, as has been done in cross-phase modulation (XPM) imaging of biological samples [XPMSS_wilson] and is frequently employed in ISRS spectroscopy and imaging [bartels2021low, Smith_IRS_Bartels_2021].
+Various techniques have been developed to measure through its modulation of probe phase in limited imaging applications measuring only the instantaneous (non-transient) cross-phase modulation (XPM) part of ) and in non-imaging applications.
+The simplest measurement detects spectral shifting proportional to the time derivative of the modulation, as has been done in XPM imaging of biological samples [29] and is frequently employed in ISRS spectroscopy and imaging [16, 26].
 A more direct measurement of probe phase can be done by interference with a reference pulse.
-Early work in ISRS focused on spectral interferometry, where the time-delayed probe and reference pulses produce interference fringes on a spectrometer [SagnacInterferometer_Kobayashi_1995, MichelsonInterferometer_Wahlstrand_2023].
+Early work in ISRS focused on spectral interferometry, where the time-delayed probe and reference pulses produce interference fringes on a spectrometer [30, 24].
 The fringe spacing is inversely proportional to probe-reference delay, and pump-induced causes the fringes to shift.
-Alternatively, probe and reference pulses can be re-timed and overlapped on a single-element detector [van2005detection, Orrit, Guillet_2015].
-Such measurements are spoiled by any relative phase instability between probe and reference, and high stability is achieved in geometries where probe and reference pulses trace a common path from the point where they are split from a common initial pulse to the point where their interference is measured after interaction with the sample.
+Alternatively, probe and reference pulses can be re-timed and overlapped on a single-element detector [31, 32, 33].
+Such measurements are spoiled by any relative phase instability between probe and reference, and require geometries where probe and reference trace a common path from the point where they are split from a common initial pulse to the point where their interference is measured after interaction with the sample.
 The challenge in common-path interferometry is to ensure that the probe interacts with the pump-induced perturbations to the sample, while the reference does not.
-This can be done in a Sagnac interferometer in which counter-propagating probe and reference pulses arrive at different times in the sample [SagnacInterferometer_Kobayashi_1995], or by propagating probe and reference along slightly different paths so that the pump overlaps spatially only with the probe [MichelsonInterferometer_Wahlstrand_2023], or by using group delay differences in bulk birefringent media to create a probe-reference pair, which arrive at different times in the sample [Schlup].
-(It should be noted that for short probe-reference delays, these methods measure a finite difference rather than directly [Wilson_SyntheticAperture_2008, bartels2021low].) Of these, the birefringent common path interferometer provides a simple means of re-timing probe and reference, enabling measurement with a single-element detector rather than a spectrometer, and has been employed for microscopic spectroscopy.
-In a transmission setup with a pair of pulse-splitting and re-timing calcites, [Orrit] measured both the real and imaginary components of the perturbed optical transmission coefficient for gold nanoparticles.
-Similarly, [Guillet_2015] made complex measurements of the perturbed optical reflection coefficient for a tungsten thin film.
-In their setup, the probe and reference are split on their way to the sample by a birefringent crystal, then after reflection from the sample re-timed by propagating in the reverse direction through the same crystal.
+This can be done in a Sagnac interferometer in which counter-propagating probe and reference pulses arrive at different times in the sample [30, 34], or by propagating probe and reference along slightly different paths so that the pump overlaps spatially only with the probe [24], or by using group delay differences in bulk birefringent media to create a probe-reference pair, which arrive at different times in the sample [35].
+(It should be noted that for short probe-reference delays, these methods measure a finite difference rather than directly [21, 16].)
+Of these, the birefringent common path interferometer provides a simple means of re-timing probe and reference, enabling measurement with a single-element detector rather than a spectrometer, and has been employed for non-imaging spectroscopy.
+In a transmission setup with a pair of pulse-splitting and re-timing calcites [32], measured both the real and imaginary components of the perturbed optical transmission coefficient for gold nanoparticles.
+Similarly, [33] made complex measurements of the perturbed optical reflection coefficient for a tungsten thin film.
+In their setup, the probe and reference are split on their way to the sample by a birefringent crystal, then after reflection from the sample, re-timed by propagating in the reverse direction through the same crystal.
 A half-wave plate can shift the measurement from phase to amplitude in both studies, permitting measurement of the complex .
 This approach to TM has yet to be implemented on a galvanometer-scanning microscope, and thus has not been applied to imaging of biological samples.
-In addition, the placement of the pump-combining dichroic before the birefringent splitter precludes rotation of pump polarization angle, which can have a large effect on time-resolved spectroscopy of biological pigments such as melanin [Grass_PolarizationMelaninPP_Warren_2022].
-Here, we extend TM with birefringent inline interferometry to a galvanometer-scanning microscope, measuring biological samples in a transmission geometry.
-In contrast to previous works [Orrit, Guillet_2015], we place the pump-combining dichroic after the probe-reference birefringent splitter, allowing for the freedom to arbitrarily rotate the pump polarization.
-We evaluate the impact that galvanometer scanning has on probe-reference re-timing and field of view, in addition to the signal-to-noise impact of polarization effects from placing the dichroic after the birefringent probe-reference splitter [DichroicAberrations].
+In addition, the placement of the pump-combining dichroic before the birefringent splitter precludes rotation of pump polarization angle, which can have a large effect on time-resolved spectroscopy of biological pigments such as melanin [36].
+Here, we extend TM with birefringent inline interferometry to a galvanometer-scanning microscope, imaging biological and crystalline samples in a transmission geometry.
+With this setup we are able to compare two primary differences between TAM and TM: (1) differences in the SNR for a particular sample at our set pump and probing wavelengths and (2) additional sample information revealed by instantaneous phase interactions (i.e.
+XPM) versus instantaneous absorption interactions (i.e.
+TPA), which is particularly revealing in an imaging geometry.
+In contrast to previous works [32, 33], we place the pump-combining dichroic after the probe-reference birefringent splitter, allowing for the freedom to arbitrarily rotate the pump polarization.
+We evaluate the impact that galvanometer scanning has on probe-reference re-timing and field of view, in addition to the signal-to-noise impact of polarization effects from placing the dichroic after the birefringent probe-reference splitter [37].
 Furthermore, we implement a balanced detection scheme that doubles the TM signal while canceling relative intensity noise from the laser.
-We demonstrate imaging applications on graphene and red blood cells (using two-photon absorption for pumping), comparing the relative signal-to-noise ratio (SNR) achievable between TAM and TM.
-Our results demonstrate the feasibility of galvanometer-scan imaging using transient phase detection and suggest that the ability to switch between TAM and TM can yield an improved SNR.
+We demonstrate spectroscopy and imaging measurements on graphene (in which TAM yields a stronger signal) and hemoglobin and red blood cells (in which TM yields a stronger signal).
+Our results demonstrate the feasibility of galvanometer-scan imaging using transient phase detection and suggest that the ability to switch between TAM and TM can yield an improved SNR for a given set of pump and probe wavelengths.
 2 Methods
-This research presents a comparison of time-resolved transient absorption versus time-resolved transient phase measurements.
-We utilize a pair of birefringent calcite crystals for separation and re-timing of the probe and reference pulses, then measure the temporal interference of the probe-reference pair following the pump-sample interaction.
 2.1 Experimental Setup
-Our laser source is generated by a Ytterbium-doped fiber laser oscillator (Menlo YLMO) which seeds a fiber chirped pulse amplifier that provides pulses at centered at with a bandwidth of nm.
-This output is split into a pump pulse train for the experiments, and another portion feeds a noncollinear optical parametric amplifier to generate the probe and reference pulses at with a bandwidth of nm.
+A Yb-doped fiber laser oscillator (Menlo YLMO) seeds a fiber chirped pulse amplifier (KMLabs) that provides pulses at centered at a fixed with a bandwidth of nm.
+This output is split into a pump pulse train for the experiments, and another portion feeds a noncollinear optical parametric amplifier (KMLabs) to generate the probe and reference pulses at with a bandwidth of nm.
 Afterwards, dispersion compensation is managed on the pump arm with a grating compressor and a prism compressor, and on the probe arm with a prism compressor.
+We used the fixed output for the pump and the tunable output for the probe to enable future studies on probe wavelength dependence.
 The system diagram after the laser source is shown in Fig.
 1.
-The probe arm is directed through delay stage which controls the delay between the pump and probe pulse trains.
+A motorized stage delays the probe relative to the pump by .
 A set of polarization optics (HWP 1, POL 1, HWP 2) cleans up the probe polarization and rotates it to with respect to the eigenaxes of a 2 mm thick calcite (CAL 1), with its optic axis oriented in the plane of the table.
 This results in an equal projection of the probe pulse along the fast (extraordinary) and slow (ordinary) axes of the crystal, respectively, resulting in two output pulses.
-The reference pulse exits first and is p-polarized followed by a delayed probe pulse that is s-polarized (see Supplementary 1A).
-The pump arm is directed through an acousto-optic modulator (AOM), which applies a 500 kHz sinusoidal amplitude modulation.
+The p-polarized reference exits first and is followed later by an s-polarized probe (see Supplementary 1A).
+The pump acquires a 500 kHz sinusoidal amplitude modulation by an acousto-optic modulator (AOM).
 A polarizer and half-wave plate provides rotatable pump polarization, before combining with the probe-reference pair in the dichroic mirror (DM) (Thorlabs DMLP900).
-The combined pulses are directed into a laser scanning microscope, with a pair of galvanometernometer scan mirrors conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm.
-A pump reject filter eliminates the pump pulse train after the condenser and the reference-probe pulse pair is relayed from the condenser to the re-timing calcite (CAL 2).
+The combined pulses are directed into a laser scanning microscope, with a pair of galvanometer scan mirrors conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm.
+The objective has a numerical aperture (NA) of 0.75, leading to an estimated diameter of for the pump, for the probe, and for the product of the two beams at the focus (neglecting aberrations and assuming Gaussian profiles).
+A pump reject filter eliminates the pump pulse train after the condenser and the reference-probe pair is relayed from the condenser to the re-timing calcite (CAL 2).
 CAL 2 is oriented so as to undo the relative delay between the probe and reference, with its slow (ordinary) axis aligned with the advanced reference pulse and fast (extraordinary) axis aligned with the delayed probe pulse.
-At the exit of the re-timing calcite, probe and reference pulses are temporally overlapped and in quadrature ( relative offset phase, adjusted by fine-tuning the azimuthal angle of the splitter calcite) but orthogonally polarized.
-A half-wave plate then rotates probe and reference so they are at with respect to the s- and p- polarization axes of a polarizing beam splitter (PBS).
-Detector A (Thorlabs PDA36A2) measures p-polarized light transmitted through the PBS and Detector B (Thorlabs PDA36A2) measures s-polarized light reflected from the PBS.
-In this configuration, an advance in probe phase relative to the pump increases A while decreasing B.
-By rotating the half-wave plate, detection can be reconfigured for TAM, rather than TM, as detailed below.
+Afterwards, probe and reference are temporally overlapped and in quadrature ( relative offset phase, adjusted by fine-tuning the azimuthal angle of the splitter calcite) but orthogonally polarized.
+A half-wave plate then rotates probe and reference so each is with respect to the axis of a polarizing beam splitter (PBS).
+Detector A (Thorlabs PDA36A2) measures the p-polarized projection and Detector B (Thorlabs PDA36A2) the s-polarized projection.
+In this configuration, an advance in probe phase relative to the reference increases A while decreasing B.
+Rotating the half-wave plate can reconfigured for TAM, rather than TM, as detailed below.
 Throughout, relay lenses are used to image the Fourier plane (collection objective back aperture) to the re-timing calcite and the detectors, minimizing the effects of lateral movement during image scanning applications.
@@
 A diagram focusing on the polarization optics is shown in Figure 2.
-The probe beam proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1).
-The polarizer ensures that we have linearly polarized light at step 1 which is p-polarized (parallel to the plane of the optics table).
-HWP 2 is used to rotate the linearly polarized light to 45 degrees (step 2).
-CAL 1 is a 2 mm thick piece of calcite cut so that the optic axis runs along the front face of the crystal (a-cut).
-It is oriented so that the extraordinary axis coincides with p-polarized light and the ordinary axis coincides with s-polarized light.
-When the probe beam is incident on CAL 1 at 45 degrees it is projected onto the ordinary () and extraordinary axes ().
+The probe proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1), resulting in linear p-polarized (parallel to the plane of the table) at step 1.
+HWP 2 rotates the polarization by (step 2).
+CAL 1 is a 2 mm thick piece of a-cut calcite (optic axis runs along the front face).
+It is oriented so that p-polarization sees the extraordinary axis and s-polarization sees the ordinary axis.
+Thus the probe is projected onto the ordinary ( [38]) and extraordinary axes ( [38]).
 Here and for are the refractive index and group index, evaluated at the center frequency of the probe pulse, .
-Two pulses exit CAL 1 (step 3) orthogonally polarized; the reference pulse is advanced and p-polarized, the probe pulse is delayed and s-polarized.
-The probe pulse pair proceeds through the microscope before encountering CAL 2.
+Two pulses thus exit CAL 1 (step 3) orthogonally polarized; the reference is advanced and p-polarized, the probe is delayed and s-polarized.
+The pair proceeds through the microscope before encountering CAL 2.
 CAL 2 is oriented opposite to CAL 1; the ordinary (slow) axis is rotated to coincide with the advanced reference pulse and the extraordinary (fast) axis coincides with the delayed probe pulse.
 The result is a re-timing of the two pulses (step 4).
-Note that the axes of CAL 2 may not align with s- and p-polarization as referenced to the table surface, but is instead based on the axes of the reference and probe polarizations after going through the microscope.
-HWP 3 is used to rotate the re-timed reference and probe polarizations such that they are 45 degrees to the s- and p- axes of the polarizing beam splitter (PBS), so that both the reference and probe pulses send a copy of themselves to DET A and DET B (step 5).
-The pulse pair copy on DET B (step 6) will be out of phase from the pulse pair on DET A (step 7).
+Note that the axes of CAL 2 may not align with s- and p-polarization as referenced to the table surface, but is instead based on the axes of the reference and probe polarizations after going through the microscope (this is due to the orientation of the detection arm which utilizes a vertical beam path relative to the tabletop).
+HWP 3 is used to rotate the re-timed reference and probe polarizations such that they are 45 degrees to the s- and p- axes of the polarizing beam splitter (PBS), so that both the reference and probe pulses arrive at both DET A and DET B (step 5).
+The pulse pair on DET B (step 6) will be out of phase from the pulse pair on DET A (step 7).
 Thus, if one detector measures destructive interference between the reference-probe pair then the other detector will measure constructive interference.
-With each sample, the initial offset phase () between the probe and reference pulses is set to a balanced position such that , where is the voltage on DET A and is the voltage on DET B.
+With each sample, the initial offset phase () between the probe and reference is set to a balanced position such that , where is the voltage on DET A and is the voltage on DET B.
 The offset phase can be set by adjusting the angle-of-incidence () of CAL 1 with p-polarized light.
@@
 The preamplifier (SR 560) measures and applies a voltage gain before the difference signal is sent into a lock-in amplifier (Liquid Instruments Moku:Lab).
-The lock-in amplifier measures amplitude changes of that coincide with the 500 kHz modulation frequency of the pump train; these measurements correspond to pump-induced phase changes .
+The lock-in amplifier measures amplitude changes of that coincide with the 500 kHz pump modulation frequency; these measurements correspond to pump-induced phase changes .
 The balanced detection signal is given by
@@
 In the balanced configuration, , and assuming that the phase perturbations resulting from are small,
-One advantage of this detection configuration is the ability to switch between measuring transient absorption () and transient phase () simply by changing the projection of the probe and reference pulses on the PBS axes [Orrit, Guillet_2015].
+One advantage of this detection configuration is the ability to switch between measuring transient absorption () and transient phase () simply by changing the projection of the probe and reference pulses on the PBS axes [32, 33].
 This projection angle is controlled by HWP 3; Figure 4 shows three projection scenarios of the reference-probe pair polarization axes onto the PBS.
@@
 3.1 Transient Phase and Absorption Measurements
-This section presents the results of measuring transient absorption and transient phase in glass, a graphene monolayer, bovine hemoglobin in solution and in live human red blood cells.
-The probe wavelength of 800 nm is considered off-resonance for the hemoglobin and red blood cell samples.
+This section presents the results of measuring transient absorption and transient phase in glass, a graphene monolayer, bovine hemoglobin in solution and fresh human red blood cells.
 3.1.1 Cross-Phase Modulation Measurement in Glass
@@
 Glass has a minimal absorption at the probe wavelength of 800 nm and the interaction that occurs between the sample and the pump pulse is instantaneous.
-Thus, the width of the XPM pulses gives us an estimate of pulse duration for our reference and probe pulses as they overlap with the pump pulse in time; these pulses are 350 fs in duration.
+When making an absorption measurement, we would expect to see a two-photon absorption peak when the pump and probe pulses were overlapped if the sample absorbed the pump wavelength.
+For a phase measurement, we are able to see cross-phase modulation (XPM) from the instantaneous optical Kerr effect, in which pump intensity changes the refractive index seen by the probe [21].
+Consequently, due to the instantaneous nature of the interaction, the width of the XPM pulses indicates a cross-correlation between pump and probe.
+(Frequency resolved optical gating (FROG) measurements were previously made that showed the probe pulse width was the limiting factor).
 3.1.2 Graphene Monolayer
-Next we explored the 2D and 3D nonlinear response of a graphene monolayer on a 0.5 mm quartz substrate placed on top of a 1 mm glass microscope slide.
-Previous works looking at graphene have utilized transient absorption to analyze graphene carrier dynamics [GrapheneTAS_SingleLayerHyperspectral_Gesuele_2019, MaterialSciApp_3_Ma_2022].
-It is a material that shows promise for use in various applications including electronics and drug delivery, making it a popular research target [Graphene_TAM_InBlood_Yang_2015].
-Here, graphene serves as a control material to compare TAM with TM since it has a strong transient absorption response at our pump and probe wavelengths.
+Next we examined a graphene monolayer on a 0.5 mm quartz substrate placed on top of a 1 mm glass microscope slide.
+Previous works looking at graphene have utilized transient absorption to analyze graphene carrier dynamics [39, 12].
+Near-infrared pumping (as in our setup) results in transient heating of free carriers [40].
+It is a material that shows promise for use in various applications including electronics and drug delivery, making it a popular research target [41].
 Time-Domain Spectroscopy
-An initial delay scan taken in the center of the sample revealed a stronger transient absorption versus transient phase response with a decay time of 2 picoseconds.
+Figure 6 shows a delay scan taken in the center of the sample.
+A stronger transient absorption versus transient phase response is observed, with a decay time of 2 ps (found by fitting to an exponential function).
 The first peak at ps corresponds to the pump-reference overlap and the second peak at ps corresponds to the pump-probe overlap.
 Imaging
-Image stacks were acquired at the edge of the graphene sample where we expected to see some variation in the deposition of graphene on the quartz substrate.
+Image stacks were acquired at the edge of the graphene sample where spatially-varying structure is expected.
+Images of pixels were acquired with 0.05 second pixel dwell time (8.3 minutes/frame).
 A probe transmissivity image was acquired after setting the half-waveplate for phase measurements and the angle of incidence for the first calcite for a "balanced" state (i.e.
@@
 From Figure 7 the edge of the sample can be seen and gives context to interpreting the image stack in Figure 8.
-An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared with output from the lock-in amplifier (V) (Figure 8).
-The transient absorption measurements still provide better SNR, but similar structural information can be seen in both techniques.
+An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared (Fig.
+8).
+Transient absorption still provides better SNR, but similar structural information can be seen in both.
 3.1.3 Hemoglobin and Red Blood Cells
@@
 First, we looked at the spectroscopy of a hemoglobin sample prepared with 9 mM of bovine hemoglobin lyophilized powder (Sigma-Aldrich H2500) in phosphate buffer solution.
-Then we imaged a droplet of whole blood placed on a poly-L-lysine coated microscope slide with a glass coverslip.
+Second, we imaged a droplet of whole blood placed on a poly-L-lysine coated microscope slide with a glass coverslip.
 The whole blood used was acquired from healthy donors by venipuncture by Dr.
@@
 Time-Domain Spectroscopy of Hemoglobin
-The hemoglobin sample showed a decay curve (top row in Figure 9).
-Interestingly, hemoglobin demonstrated a pump polarization-dependent signal with differences between the polarization sensitivity of the transient absorption versus transient phase measurements (bottom row of Figure 9).
-This would be impossible to observe if the pump-combining dichroic was placed before the probe-reference splitting calcite like in [Orrit, Guillet_2015].
+The hemoglobin sample showed a decay curve (top row, Fig.
+9).
+For a discussion of the physics behind the transient absorption response of hemoglobin at these wavelengths, we refer the reader to [6].
+Interestingly, hemoglobin demonstrated a pump polarization-dependent signal with differences between the polarization sensitivity of the transient absorption versus transient phase measurements (bottom row, Fig.
+9).
+This would be impossible to observe if the pump-combining dichroic was placed before the probe-reference splitting calcite like in [32, 33].
 With transient absorption the pump-reference overlap showed a polarization dependence while the pump-probe overlap did not.
@@
 With transient phase both overlaps demonstrated a polarization dependence; the pump-reference overlap shows the largest signal for the s-polarized pump and the pump-probe overlap shows the largest signal for the p-polarized pump.
+The polarization dependence is stronger at shorter delays (< 1 ps), similar to pump-probe anisotropy measurements on other hemeproteins [42].
 Red Blood Cell Imaging
-Imaging human red blood cells, we expected to see a predominant signal from hemoglobin and a similar SNR from the two techniques (based on the results from the delay scans).
-However, the transient phase demonstrated a stronger signal than transient absorption in this case with clearer structural details due to XPM.
-Figure 10 shows a comparison of TM and TAM for imaging the transient decay in red blood cells.
+Imaging human red blood cells, we expected to see a predominant signal from hemoglobin and a similar signal level from the two techniques (based on the results from the delay scans).
+However, TM demonstrated a much larger signal, with clearer structural details (Fig.
+10).
 All of the images are shown on the same intensity scale; the overlaps with the pump-reference at ps and with the pump-probe at ps show the strongest signals and good structural detail for TM.
@@
 4.1 Impacts of Dichroic Beam-Combiner on SNR
-The pump-combining dichroic was placed after the probe-reference splitting calcite, unlike prior works which placed the dichroic before the splitter to avoid splitting the pump [Orrit, Guillet_2015].
+The pump-combining dichroic was placed after the probe-reference splitting calcite, unlike prior works which placed the dichroic before the splitter to avoid splitting the pump [32, 33].
 This enabled characterization of polarization dependence, as well as selection of relative pump-probe polarization that maximizes signal for imaging.
-It should be noted, however, that the dielectric coating of the dichroic is subject to polarization aberrations, including dichroism (polarization-dependent absorption), depolarization, dispersion, and retardance [DichroicAberrations].
+It should be noted, however, that the dielectric coating of the dichroic is subject to polarization aberrations, including dichroism (polarization-dependent absorption), depolarization, dispersion, and retardance [37].
 These effects can depend on both incident polarization state and angle of incidence, and can ultimately interfere with our detection scheme.
@@
 4.2 Impact of Galvanometer-scanning on Re-timing Phase in Imaging
-Galvanometer-scanning microscopy is a new development in TM and comes at the price of an angle-dependent re-timing phase present across the scanned field of view.
-In order to determine how great an impact galvanometer-dependent re-timing phase has on this system, a cross-phase modulation (XPM) image was taken in a glass slide at different fields of view.
-Since glass is homogeneous and has no transient signal, a flat signal would be expected across the entire field of view in the absence of galvanometer-dependent re-timing phase.
+One drawback to this TM approach is an angle-dependent re-timing phase offset across the scanned field of view (FOV).
+To characterize this, an XPM image was taken in a glass slide at different FOV.
+Since glass is homogeneous and has no transient signal, a flat signal would be expected across the entire FOV in the absence of spatially-dependent re-timing.
 The measured data shown in Figure 11 indicates a phase shift that depends on scan angle.
-The images shown in Figure 11 have had the on-axis phase shift DC subtracted so only the spatially dependent phase shift is portrayed.
-Clearly, the larger field of view sees a stronger galvanometer-dependent phase shift .
-Overall, the signal measured across a field of view will equal the sum + .
-Decreasing the field of view or increasing the magnification of the beam at the re-timing calcite (CAL 2) could help minimize the re-timing phase effects (Supplementary 4).
+The images shown in Figure 11 have had the on-axis (center of image) phase shift subtracted so only the spatially dependent phase shift is portrayed.
+Clearly, the larger FOV sees a stronger galvanometer-dependent phase shift .
+Overall, the signal measured across a FOV will equal the sum + .
+Decreasing the FOV or increasing the magnification of the beam at the re-timing calcite (CAL 2) could help minimize the re-timing phase effects (Supplementary 4).
 In order to completely eliminate this effect, de-scanning could be added in prior to CAL 2.
 4.3 Trade-offs
-One significant benefit of measuring TM is that the instantaneous cross-phase modulation (XPM) is measured in addition to the transient phase signal.
-The XPM contributes additional context about the structure of the sample related to nonlinear refractive index variations across the sample [XPMSS_wilson].
-TAM will also measure instantaneous phenomena, such as two-photon absorption, but this does not give the same structural clarity as XPM and is instead specific only to materials and molecules resonant with the sum of the pump and probe energies.
+One significant benefit of measuring TM is that XPM is measured in addition to pump-induced transients.
+The XPM contributes additional context about the structure of the sample related to nonlinear refractive index variations across the sample [29], though it carries no information about excited-state lifetimes.
+TAM will also measure instantaneous phenomena, such as two-photon absorption, but this does not give the same structural clarity as XPM and is instead specific only to materials and molecules resonant with the sum of the pump and probe energies (with peak intensities sufficient for exciting nonlinear processes).
 The microscope design presented here allows for a simple conversion between transient phase and transient absorption measurements.
@@
 This was discussed above for the case of the dichroic beam splitter, but the effects of birefringent samples can also be considered.
-For thin, weakly birefringent samples the polarization relationship of the reference and probe pulses is at risk of changing, and we would expect the effects to be similar to the dichroic in reducing SNR.
+Another potential drawback to this setup is the need for polarization purity between the probe and reference.
+Any loss in the ability to separate probe and reference could degrade SNR, e.g.
+from the dichroic, microscope optics, and the sample.
 Thick birefringent samples run the risk of not only changing the polarization relationship of the reference and probe pulses but also of additional pulse splitting or temporal separation.
@@
 5 Conclusion
-The purpose of this research was to extend TM to a galvanometer-scanning configuration, as a complement to TAM for biological samples.
-In sum, the scanning galvanometer TM approach described here can offer advantages over TAM, such as XPM structural contrast and potentially higher signal levels, but is susceptible to degradation from polarization aberrations in the optical setup and sample birefringence, and has more limited field of view.
+The purpose of this research was to extend TM, a measurement of pump-induced refractive index changes measured with interferometric probe and reference pulses, to a galvanometer-scanning configuration as a complement to TAM for biological samples.
+In sum, TM can offer advantages over TAM, such as XPM structural contrast and potentially higher signal levels (as was the case here for hemoglobin and red blood cells), but is susceptible to degradation from polarization aberrations in the optical setup and sample birefringence, and has more limited FOV.
 Thus, it is beneficial to be able to acquire both TAM and TM with a simple rotation of a waveplate, as described here.
-In the future two changes to the microscope design here could be made to improve the utility for TM.
+In the future, two changes could be made to improve the utility for TM.
 Firstly, the dichroic beam combiner could be swapped for a 50/50 beamsplitter.
@@
 Secondly, de-scanning could be added prior to the re-timing calcite to remove the spatial dependence on re-timing phase.
-Overall, TM shows promise as a complementary technique to TAM, particularly for setups that may be limited in their ability to vary probe wavelengths.
+Overall, TM shows promise as a complementary technique to TAM, particularly for setups that may be limited in their ability to vary probe wavelengths and as a way to enhance structural context of a transient image through its simultaneous measurement of XPM.
 Funding This research was supported by the National Institute of General Medical Sciences of the National Institutes of Health under award number R21GM135772 and by the Biophotonics Program of the National Science Foundation under award number 1943595.

### v3

@@
 With this setup we are able to compare two primary differences between TAM and TM: (1) differences in the SNR for a particular sample at our set pump and probing wavelengths and (2) additional sample information revealed by instantaneous phase interactions (i.e.
-XPM) versus instantaneous absorption interactions (i.e.
-TPA), which is particularly revealing in an imaging geometry.
+cross-phase modulation, XPM) versus instantaneous absorption interactions (i.e.
+two-photon absorption, TPA), which is particularly revealing in an imaging geometry.
 In contrast to previous works [32, 33], we place the pump-combining dichroic after the probe-reference birefringent splitter, allowing for the freedom to arbitrarily rotate the pump polarization.
@@
 1.
-A motorized stage delays the probe relative to the pump by .
+A motorized stage delays the probe beam (prior to splitting into a reference-probe pulse-pair) relative to the pump by .
+(After the probe is split by CAL 1 into a reference-probe pulse-pair this temporal separation, is defined between the pump and the leading pulse from the pulse-pair, i.e.
+the reference pulse, inset B).
 A set of polarization optics (HWP 1, POL 1, HWP 2) cleans up the probe polarization and rotates it to with respect to the eigenaxes of a 2 mm thick calcite (CAL 1), with its optic axis oriented in the plane of the table.
@@
 The p-polarized reference exits first and is followed later by an s-polarized probe (see Supplementary 1A).
-The pump acquires a 500 kHz sinusoidal amplitude modulation by an acousto-optic modulator (AOM).
-A polarizer and half-wave plate provides rotatable pump polarization, before combining with the probe-reference pair in the dichroic mirror (DM) (Thorlabs DMLP900).
-The combined pulses are directed into a laser scanning microscope, with a pair of galvanometer scan mirrors conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm.
+The pump acquires a 500 kHz sinusoidal amplitude modulation by an acousto-optic modulator (AOM; TEM-100-10-1050, Brimrose Corp., Baltimore, MD).
+A polarizer and half-wave plate provide rotatable pump polarization, before combining with the probe-reference pair in the dichroic mirror (DM) (Thorlabs DMLP900).
+The combined pulses are directed into a laser scanning microscope, with a pair of galvanometer scan mirrors (QS-7, Nutfield Technology, Inc., Windham, NH) conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm.
 The objective has a numerical aperture (NA) of 0.75, leading to an estimated diameter of for the pump, for the probe, and for the product of the two beams at the focus (neglecting aberrations and assuming Gaussian profiles).
@@
 A diagram focusing on the polarization optics is shown in Figure 2.
-The probe proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1), resulting in linear p-polarized (parallel to the plane of the table) at step 1.
+The probe proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1), resulting in linear p-polarization (parallel to the plane of the table) at step 1.
 HWP 2 rotates the polarization by (step 2).
@@
 Ultimately our setup is capable of measuring an excited-state phase change of from the initial offset point.
-The voltage measured on each detector is the time-integrated interference between the reference and probe; since our final signal is we end up with a 2X increase in measurable signal than if we were conducting our measurements with only one detector.
-The preamplifier (SR 560) measures and applies a voltage gain before the difference signal is sent into a lock-in amplifier (Liquid Instruments Moku:Lab).
+The voltage measured on each detector is the time-integrated interference between the reference and probe; since our final signal is we end up with a 2X increase in measurable signal than if we were conducting our measurements with a single detector.
+(It should be noted, though, than both TAM and TM can be done with only one detector.) The preamplifier (SR 560) measures and applies a voltage gain before the difference signal is sent into a lock-in amplifier (Liquid Instruments Moku:Lab).
 The lock-in amplifier measures amplitude changes of that coincide with the 500 kHz pump modulation frequency; these measurements correspond to pump-induced phase changes .
@@
 ) with the pump blocked.
-From Figure 7 the edge of the sample can be seen and gives context to interpreting the image stack in Figure 8.
+From Figure 7 the edge of the sample can be seen in this transmissivity image (left-most image) and gives context to interpreting the image stack in Figure 8.
+Within the transmissivity image the edge of the sample is seen at the transition from light to dark (where the light pixels correspond to the sample)
 An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared (Fig.
 8).
-Transient absorption still provides better SNR, but similar structural information can be seen in both.
+For this sample, we see that transient absorption provides a stronger signal at overlap with both the probe and reference (-1.20 ps and 0.00 ps) than transient phase.
 3.1.3 Hemoglobin and Red Blood Cells